OpenStudy (anonymous):

What is the measure of an arc that is intercepted by an inscribed angle of 30°?

OpenStudy (anonymous):

i wish i knew what this meant

OpenStudy (anonymous):

*pounds brain to get back info from pre-calc 2 years ago*

OpenStudy (anonymous):

The rule for inscribed angles is 2(theta): http://en.wikipedia.org/wiki/Inscribed_angle Since we have 30 degrees the central angle is 60 degrees. 60/360 = 1/6 of the circumference. Circumference = 2pi r The arc = (1/6) 2 pi r = (pi)(r)/3

OpenStudy (anonymous):

An inscribed angle is an angle with its vertex on a circle and with sides that contain chords of the circle.

OpenStudy (anonymous):

The measure of an inscribed angle is equal to one-half the degree measure of its intercepted arc.

OpenStudy (amistre64):

correct. if this means something akin to this, then the arc is equal to twice the angle.